In each of the independent cases below, indicate the taxable portion and return of capital, assume sufficient basis for each distribution.

accumulated E&P

Beginning of year

current E&P Year end cash distribution taxable amount return of capital
a. $40,000 $20,000 $60,000 ? ?
b. 40,000 20,000 70,000 ? ?
c. (40,000) 40,000 30,000 ? ?
d. (10,000) (20,000) 15,000 ? ?
e. 40,000 (15,000) 35,000 ? ?
Answer #1

a):-Taxable income =$(40000-20000)=$20000,Return on capital=$20000/60000=1/3x100=33.33%

b):-Taxable income=$(40000-20000)=$20000,Return on capital=$20000/70000=2/7x100=28.57%

c):-Taxable income=$(-40000-40000)=$(80000) ,i,e NIL,,Return on capital=$-80000/30000=8/3x100=266%Loss

d):-Taxable income=$(-10000+20000)=$10000,Return on capital=$10000/15000=2/3x100=66.67%

e):-Taxable income=$(40000+15000)=$55000,Return on capital=$55000/35000=1.67x100=167%